What is the maximum and minimum of f(x) = 20sin3(pi)x + 140 where 0
Answer by stanbon(75887)

What is the maximum and minimum of f(x) = 20sin3(pi)x + 140 where 0 ----
amplitude = 20
vertical shift is 140
----
max value = 140+10 = 150
min value = 140-10 = 130
=============================
Cheers,
Stan H.